OpenStudy (anonymous):

Prove that (tanX + sinX)(1-cosX) = sin^2XtanX

OpenStudy (loser66):

expand the left hand side, what do you have?

OpenStudy (anonymous):

tanX - tanXcosX + sinX - sinXcosX

OpenStudy (loser66):

tanx cos x = sinx ,so it cancel out with + sin x ,right? so far, we have tan x - sinx cosx

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

hey, recheck the right hand side, is it right? is it \(sin^2 x tanx\)

OpenStudy (anonymous):

Yes that's correct

OpenStudy (loser66):

ok, let me try, hehehe. I jump into the problem, not sure I can get the answer or not,

OpenStudy (anonymous):

Haha okay thank you

OpenStudy (loser66):

ok, factor tan out

OpenStudy (loser66):

you get the answer, hehehe... I am lucky. !!!

OpenStudy (science0229):

Sorry to interrupt, but the left side can't be sin^2XtanX... It's supposed to be sin^3XtanX...

OpenStudy (loser66):

the second term is sinx cosx = \(\dfrac{sinx cos x*cosx}{cosx}= tan x cos^2x\)

OpenStudy (anonymous):

The left side is not sin^2XtanX, the right side is. @Loser66, what do you mean factor tan out?

OpenStudy (loser66):

now back to it tan x - tanx cos^2 x= tan x(1-sin^2x) = tan x sin^2x , oh yea!! this is right

OpenStudy (anonymous):

I think I am lost...

OpenStudy (loser66):

ok, we have tanx - sinx cosx, right?

OpenStudy (raden):

alternative : see that : tan(x) + sin(x) = tan(x) (1 + cos(x)) so, (tanX + sinX)(1-cosX) = tan(x) (1 + cos(x)) (1-cos(x)) = tan(x) (1 - cos^2(x)) = tan(x) sin^2 (x)

OpenStudy (loser66):